I think I'm missing a step in my answer to this problem, did I simplify this right? help?
84 - 26 x 3 + 98 / (-7)
=84-78-14
= 84-92= -8

Answer by Alan3354(30924)

84 - 26 x 3 + 98 / (-7)
=84-78-14
= 84-92= -8
-------------
Looks right, if only the last term, the 98, is the numerator.
It might be
(84 - 26 * 3 + 98) / (-7)
= 104/-7 = -104/7
No way to know without parentheses.